A shipment of 100 bags of sugar is rumoured to have some bags that contain illegal drugs. Local customs officials decide to screen shipments by randomly testing 4 bags. If any are found to contain drugs, the whole consignment is seized.

Compute the probability that the shipment consisting of 95 bags of sugar and 5 bags of illegal drugs will be detected in this manner.


1
Expert's answer
2021-05-18T13:54:50-0400

In how many ways can 4 bags be chosen from 100 bags


"\dbinom{100}{4}"

In how many ways can 4 bags of sugar be chosen from 96 bags of sugar


"\dbinom{95}{4}"

What is the probability that Local customs officials will choose 4 bags of sugar


"P(4)=\dfrac{\dbinom{95}{4}}{\dbinom{100}{4}}=\dfrac{95!4!(100-4)!}{4!(95-4)!100!}"

"=\dfrac{95(94)(93)(92)}{100(99)(98)(97)}"

The probability that the shipment consisting of 95 bags of sugar and 5 bags of illegal drugs will be detected is


"P(be\ detected)=1-\dfrac{95(94)(93)(92)}{100(99)(98)(97)}"

"\approx0.188125"